# Quantifying non-Hermiticity using single- and many-particle quantum properties

### v1

@@ III.1.5 Non-Hermiticity score of von Neumann entropy (dynamic case)
Further above, we have discussed the non-Hermiticity score for the VNE when computed for the special cases of $\rho_{RR}(t)\overset{t=\infty}{\rightarrow}\rho_{RR}^{k}=|R_{k}\rangle\langle R_{k}|$ and $\rho_{LL}(t)\overset{t=\infty}{\rightarrow}\rho_{LL}^{k}=|L_{k}\rangle\langle L_{k}|$. In this section, we consider a more general scenario where $\rho_{RR}(t)$ is obtained following Eq. (7) (and mutatis mutandis for $\rho_{LL}(t)$). Note that, in this case, the VNE is not an entanglement measure, as the global states $\rho_{RR}(t)$ and $\rho_{LL}(t)$ are not pure. In Fig. 3(d)-(f), we plot the time-dependent behavior of $SC^{\mathcal{S}}(t)=|\mathcal{S}[\rho_{RR}(t)]-\mathcal{S}[\rho_{LL}(t)]|$, for the same parameter values as in panels (a)-(c). Similar to the purity score, $SC^{\mathcal{S}}$ exhibits non-monotonic behavior with time and its maximum value computed for the considered time period, $SC^{\mathcal{S}}_{\max}=\max_{t\in[0,100]}\{SC^{\mathcal{S}}(t)\}$, again remains low for high values of $\alpha$. However, this score achieves significant values at small $\alpha$ even at large $\delta$.
@@ III.2 Non-Hermitian interacting Hatano–Nelson model
$H^{\mathrm{HN}}_{\mathrm{nh}}=\sum_{i}^{N}-J(e^{\chi}c^{\dagger}_{i}c_{i+1}+e^{-\chi}c^{\dagger}_{i+1}c_{i})+V\hat{n}_{i}\hat{n}_{i+1}\,.$ (20)
Here, the hopping strength $J=1$ scales the Hamiltonian, $c_{i}^{\dagger}$($c_{i}$) is the creation (annihilation) operator for a spinless fermionic particle at site $i$, $\hat{n}_{i}=c_{i}^{\dagger}c_{i}$ the single-mode number operator, and $V$ is the strength of nearest-neighbor interaction. Moreover, we restrict our studies to the half-filling sector and employ periodic (anti-periodic) boundary conditions for odd (even) $N/2$. The notion of non-Hermiticity is introduced by asymmetric hopping controlled by the parameter $\chi$, where the model reduces to the Hermitian version at $\chi=0$. The model has parity-time (PT) reversal symmetry, which results in eigen energies appearing as complex conjugate pairs. In Ref. [43], it is reported that as the strength of the interaction ($V$) increases, the model undergoes two PT symmetry-breaking transitions: One of them is due to an exceptional point between the two lowest excited states and marked by the non-analytic behavior of the finite-size level-spacing $\Delta_{01}:=\mathrm{Re}(\lambda_{1}-\lambda_{0})$. The other corresponds to a full collapse of the complex many-body spectrum onto the real axis in a finite-size system. The former transition is accompanied by a first-order symmetry-breaking transition from a PT broken gapless phase to a PT-symmetric charge-density phase. Similar characteristics are also reported in the hardcore bosonic version of the model with periodic boundary condition [44]. In what follows, we present the results for the degree of Hamiltonian non-Hermiticity of the model as well as single- and multi-site non-Hermiticity scores, and also contrast their trend against the above-mentioned transitions.
@@
To begin with, we examine how the asymmetric coupling parameterized by $\chi$ and interaction strength $V$ affect the degree of non-Hermiticity ($\mathcal{D}$) of the Hamiltonian $H^{\mathrm{HN}}_{\mathrm{nh}}$, plotted in Fig. 1(b). At $V=0$, $\mathcal{D}$ grows monotonously with $\chi$ and eventually saturates to the theoretical maximum of $\mathcal{D}\approx 2$. However, turning on even a small amount of interaction has a detrimental effect on $\mathcal{D}$: it attains a significant non-zero value only beyond a non-vanishing threshold coupling, $\chi_{c}$. Interestingly, we further observe that up to $V\approx 10^{3}$, $\chi_{c}$ increases almost linearly with $V$. At even larger interaction strengths, $\mathcal{D}$ remains zero up to the largest values of $\chi$ considered in our analysis ($\chi=5.4$).
@@
As these results show, interactions suppress the non-Hermiticity of the model significantly, as the distance between $H_{\mathrm{nh}}^{\mathrm{HN}}$ and $(H_{\mathrm{nh}}^{\mathrm{HN}})^{\dagger}$ becomes insignificant in comparison to the norm of $H_{\mathrm{nh}}^{\mathrm{HN}}$. Interestingly, the domain of vanishing $\mathcal{D}$ coincides with the region of non-analyticity of $\Delta_{01}$, which can be read off from Fig. 1(b) and its inset. This indicates that $\mathcal{D}$ can be useful in detecting the PT symmetry-breaking transition present in the model. In turn, this may signify the importance of the normalization of $\mathcal{D}$ considered in Eq. (13). In fact, the unnormalized incarnation of $\mathcal{D}$ does not reveal any transition (see Appendix C). In the following, we will examine to what extent such a behavior reflects in the behavior of single- and multi-site quantum properties related to the model.
@@
Figure 4: Comparison of non-Hermiticity at the level of the Hamiltonian and single- and multi-party physical properties of the non-Hermitian Hatano–Nelson model ($H_{\mathrm{nh}}^{\mathrm{HN}}$). We compute the non-Hermiticity score for both the number operator $\hat{n}$ and half-chain VNE $\mathcal{S}$ for all the right and left eigenvectors of the model and finally compute the maximum among all of them. (a) Depicts the behavior of $SC^{\hat{n}}_{\max}$ in the $V-\chi$ plane. We note that for a large region in the parameter space, the behavior of $SC^{\hat{n}}_{\max}$ remains almost complementary to $\mathcal{D}$, as shown in Fig. 1(b). A similar plot for the half-chain VNE is presented in (b), where for the same set of parameters, we plot the behavior of $SC^{\mathcal{S}}_{\max}$. We note that $SC^{\mathcal{S}}_{\max}$ behaves differently from both $SC^{\hat{n}}_{\max}$ as well as the degree of non-Hermiticity ($\mathcal{D}$). In particular, $SC^{\mathcal{S}}_{\max}$ attains significant value only when both $V$ and $\chi$ take large values. However, closer inspection shows the qualitative agreement between the trend of $SC^{\hat{n}}_{\max}$ and $SC^{\mathcal{S}}_{\max}$. The white circles correspond to the PT symmetry-breaking transition marked by the non-analytic behavior of $\Delta_{01}$ defined in Sec. III.2 and shown in the inset of Fig. 1. We notice the agreement between the PT symmetry-breaking transition and domain for vanishing scores. The data is reported for $N=12$ sites, half-filling, and with anti-periodic boundary conditions.
@@ III.2.2 Single-site properties: non-Hermiticity score of number operator $\hat{n}$
Figure 5: Global characteristics of $\overline{{SC}^{\mathcal{F}}_{\mathrm{nh}}}$ for the non-Hermitian Hatano–Nelson model ($H_{\mathrm{nh}}^{\mathrm{HN}}$). We compute the total number of elements of $\overline{{SC}^{\hat{n}}_{\mathrm{nh}}}$ and $\overline{{SC}^{\mathcal{S}}_{\mathrm{nh}}}$ with the conditions ${SC}^{\hat{n}}_{\mathrm{nh}}[k]\geq\mathcal{E}_{\mathrm{Th}}^{\hat{n}}$ and ${SC}^{\mathcal{S}}_{\mathrm{nh}}[k]\geq\mathcal{E}_{\mathrm{Th}}^{\mathcal{S}}$, respectively, where $k=1,2,...,D$ with $D$ being the Hilbert space dimension. These quantities, scaled by $D$, are shown in (a) and (b). We obtain a similar trend as in Fig. 4. Data for threshold values $\mathcal{E}_{\mathrm{T}h}^{\hat{n}}=\mathcal{E}_{\mathrm{T}h}^{\mathcal{S}}=0.1$. Similar to Figs. 1 and 4, the white circles correspond to the PT symmetry-breaking transition. The data is reported for $N=12$ sites, half-filling, and with anti-periodic boundary conditions.
To study the impact of non-Hermiticity on single-site properties, we analyze the non-Hermiticity score of the number operator for the first site, $\hat{n}_{1}$. The maximum score over all eigenstates, $SC^{\hat{n}}_{\max}\equiv||\overline{{SC}^{\hat{n}}_{\mathrm{nh}}}||_{\infty}$, is plotted in Fig. 4(a). Importantly, $SC^{\hat{n}}_{\max}$ exhibits an almost complementary behavior to the non-Hermiticity of the Hamiltonian $\mathcal{D}$. In particular, at low values of interaction, $SC^{\hat{n}}_{\max}$ becomes almost zero. Once the strength of the interaction is increased, expectation values of the single-site occupation evaluated in right and left eigenvectors start to exhibit a disparity, which is captured by significantly non-zero values of $SC^{\hat{n}}_{\max}$ within an intermediate regime of $\chi$. Throughout the entire window of interaction strengths considered, $SC^{\hat{n}}$ vanishes at small $\chi$, while beyond $V\approx 10^{2}$ the non-vanishing region extends up to the maximum value of $\chi$ that we have considered. The region of $V$ where $SC^{\hat{n}}_{\max}$ becomes finite for any $\chi$ resides in close proximity to the PT symmetry-breaking transition marked by the non-analytic point of $\Delta_{01}$. Thereby, $SC^{\hat{n}}_{\max}$ can be considered as another candidate in identifying the PT symmetry-breaking transition of the model.
@@
Unlike the case of $H_{\mathrm{nh}}^{\mathrm{Bell}}$, here the spectrum of the Hamiltonian consists of a large number of eigenvectors. Therefore, to illustrate the non-Hermitian behavior across the spectrum of the model, in Fig. 5(a) we plot the quantity $\mathcal{G}^{\hat{n}}$, see Sec. II.3.2, choosing as threshold value $\mathcal{E}_{\mathrm{T}h}^{\hat{n}}=0.1$. The behavior is qualitatively similar to the one of $SC^{\hat{n}}_{\max}$, showing that in this case, the maximal score is representative of the behavior across the spectrum.
@@ IV Discussion
In this work, we have introduced a formalism that quantifies different notions of non-Hermiticity, which can emerge in any generic non-Hermitian system. In particular, as first quantifier, we have introduced the Hamiltonian non-Hermiticity, a measure that computes the normalized distance between a non-Hermitian Hamiltonian $H_{\mathrm{nh}}$ and its hermitian conjugate $(H_{\mathrm{nh}})^{\dagger}$. We have introduced a second quantifier to capture a notion of non-Hermiticity at the level of physically computable quantities related to the model. This quantifier has been motivated by noticing that in the non-Hermitian paradigm one can prepare quantum ensembles of right or left eigenvectors, which often yield incompatible results with respect to single- and multi-site quantum properties. Based on this observation, we have introduced the non-Hermiticity score that characterized the difference between observables in the right and left eigenensemble. These quantifiers enable a comparative study of the degree of non-Hermiticity across different regions of parameter space, for static as well as dynamic properties.
@@
In summary, in our work we have opened a new route to examining the quantum properties of any generic non-Hermitian system. It permits to quantify non-Hermiticity, not only along the traditional lines that solely focuses on the Hamiltonian of the model but also in respect to its influence onto observable properties. We believe our formalism can be instrumental in characterizing exotic quantum many-body phases, identifying potential non-Hermitian phase transitions, as well as in the preparation of resourceful quantum states for quantum technologies. As an application of this proposed framework, in future work we aim to apply it to gain a deeper understanding of non-Hermitian topological systems.

### v2

@@ III.1.5 Non-Hermiticity score of von Neumann entropy (dynamic case)
Further above, we have discussed the non-Hermiticity score for the VNE when computed for the special cases of $\hat{\rho}_{RR}(t)\overset{t=\infty}{\rightarrow}\hat{\rho}_{RR}^{k}=|R_{k}\rangle\langle R_{k}|$ and $\hat{\rho}_{LL}(t)\overset{t=\infty}{\rightarrow}\hat{\rho}_{LL}^{k}=|L_{k}\rangle\langle L_{k}|$. In this section, we consider a more general scenario where $\hat{\rho}_{RR}(t)$ is obtained following Eq. (14) (and mutatis mutandis for $\hat{\rho}_{LL}(t)$). Note that, in this case, the VNE is not an entanglement measure, as the global states $\hat{\rho}_{RR}(t)$ and $\hat{\rho}_{LL}(t)$ are not pure. In Fig. 3(d)-(f), we plot the time-dependent behavior of $SC^{\mathcal{S}}(t)=|\mathcal{S}[\hat{\rho}_{RR}(t)]-\mathcal{S}[\hat{\rho}_{LL}(t)]|$, for the same parameter values as in panels (a)-(c). Similar to the purity score, $SC^{\mathcal{S}}$ exhibits non-monotonic behavior with time and its maximum value computed for the considered time period, $SC^{\mathcal{S}}_{\max}=\max_{t\in[0,100]}\{SC^{\mathcal{S}}(t)\}$, again remains low for high values of $\alpha$. However, this score achieves significant values at small $\alpha$ even at large $\delta$.
@@ III.2 Non-Hermitian interacting Hatano–Nelson model
$\hat{H}^{\mathrm{HN}}_{\mathrm{nh}}=\sum_{i}^{N}-J(e^{\chi}\hat{c}^{\dagger}_{i}\hat{c}_{i+1}+e^{-\chi}\hat{c}^{\dagger}_{i+1}\hat{c}_{i})+V\hat{n}_{i}\hat{n}_{i+1}\,.$ (8)
Here, the hopping strength $J=1$ scales the Hamiltonian, $\hat{c}_{i}^{\dagger}$($\hat{c}_{i}$) is the creation (annihilation) operator for a spinless fermionic particle at site $i$, $\hat{n}_{i}=\hat{c}_{i}^{\dagger}\hat{c}_{i}$ the single-mode number operator, and $V$ is the strength of nearest-neighbor interaction. Moreover, we restrict our studies to the half-filling sector and employ periodic (anti-periodic) boundary conditions for odd (even) $N/2$. The notion of non-Hermiticity is introduced by asymmetric hopping controlled by the parameter $\chi$, where the model reduces to the Hermitian version at $\chi=0$. After explaining in Appendix F how non-Hermitian dynamics can be understood operationally as a specific case of open quantum system dynamics, we present in Appendix G a detailed discussion on how the dynamics of the Hatano-Nelson Hamiltonian—and its Hermitian conjugate—can be implemented in modern cold atom quantum systems via open quantum system techniques. The model has parity-time (PT) reversal symmetry, which results in eigen energies appearing as complex conjugate pairs. In Ref. [31], it is reported that as the strength of the interaction ($V$) increases, the model undergoes two PT symmetry-breaking transitions: One of them is due to an exceptional point between the two lowest excited states and marked by the non-analytic behavior of the finite-size level-spacing $\Delta_{01}:=\mathrm{Re}(\lambda_{1}-\lambda_{0})$. The other corresponds to a full collapse of the complex many-body spectrum onto the real axis in a finite-size system. The former transition is accompanied by a first-order symmetry-breaking transition from a PT broken gapless phase to a PT-symmetric charge-density phase. Similar characteristics are also reported in the hardcore bosonic version of the model with periodic boundary condition [32]. In what follows, we present the results for the degree of Hamiltonian non-Hermiticity of the model as well as single- and multi-site non-Hermiticity scores, and also contrast their trend against the above-mentioned transitions.
@@
To begin with, we examine how the asymmetric coupling parameterized by $\chi$ and interaction strength $V$ affect the degree of non-Hermiticity ($\mathcal{D}$) of the Hamiltonian $\hat{H}^{\mathrm{HN}}_{\mathrm{nh}}$, plotted in Fig. 1(b). Note that here, and in all subsequent plots, the parameter $V$ is presented on a logarithmic scale. At $V=0$, $\mathcal{D}$ grows monotonously with $\chi$ and eventually saturates to the theoretical maximum of $\mathcal{D}\approx 2$. However, turning on even a small amount of interaction has a detrimental effect on $\mathcal{D}$: it attains a significant non-zero value only beyond a non-vanishing threshold coupling, $\chi_{c}$. Interestingly, we further observe that up to $V\approx 10^{3}$, $\chi_{c}$ increases almost linearly with $V$. At even larger interaction strengths, $\mathcal{D}$ remains zero up to the largest values of $\chi$ considered in our analysis ($\chi=5.4$).
@@
As these results show, interactions suppress the non-Hermiticity of the model significantly, as the distance between $\hat{H}_{\mathrm{nh}}^{\mathrm{HN}}$ and $(\hat{H}_{\mathrm{nh}}^{\mathrm{HN}})^{\dagger}$ becomes insignificant in comparison to the norm of $\hat{H}_{\mathrm{nh}}^{\mathrm{HN}}$. Interestingly, the domain of vanishing $\mathcal{D}$ remains close to the region of non-analyticity of $\Delta_{01}$, which can be read off from Fig. 1(b) and its inset. This indicates that $\mathcal{D}$ can be useful in qualitatively detecting the PT symmetry-breaking transition present in this model. Furthermore, in Appendix A, we report that the Frobenius norm exhibits qualitatively similar behavior. However, it is important to note that whether this behavior holds for any generic non-Hermitian Hamiltonian remains an open question and warrants further investigation. In turn, this may signify the importance of the normalization of $\mathcal{D}$ considered in Eq. (1). In fact, the unnormalized incarnation of $\mathcal{D}$ does not reveal any transition (see Appendix B). In the following, we will examine to what extent such a behavior reflects in the behavior of single- and multi-site quantum properties related to the model.
@@
Figure 4: Comparison of non-Hermiticity at the level of the Hamiltonian and single- and multi-party physical properties of the non-Hermitian Hatano–Nelson model ($\hat{H}_{\mathrm{nh}}^{\mathrm{HN}}$). We compute the non-Hermiticty score for both the number operator $\hat{n}$ and half-chain VNE $\mathcal{S}$ for all the right and left eigenvectors of the model and finally compute the maximum among all of them. (a) Depicts the behavior of $SC^{\hat{n}}_{\max}$ in the $V-\chi$ plane. We note that for a large region in the parameter space, the behavior of $SC^{\hat{n}}_{\max}$ remains almost complementary to $\mathcal{D}$, as shown in Fig. 1(b). A similar plot for the half-chain VNE is presented in (b), where for the same set of parameters, we plot the behavior of $SC^{\mathcal{S}}_{\max}$. We note that $SC^{\mathcal{S}}_{\max}$ behaves differently from both $SC^{\hat{n}}_{\max}$ as well as the degree of non-Hermiticity ($\mathcal{D}$). In particular, $SC^{\mathcal{S}}_{\max}$ attains significant value only when both $V$ and $\chi$ take large values. However, closer inspection shows the qualitative agreement between the trend of $SC^{\hat{n}}_{\max}$ and $SC^{\mathcal{S}}_{\max}$. The white circles correspond to the PT symmetry-breaking transition marked by the non-analytic behavior of $\Delta_{01}$ defined in Sec. III.2 and shown in the inset of Fig. 1. We notice the qualitative agreement between the PT symmetry-breaking transition and domain for vanishing scores. The data is reported for $N=12$ sites, half-filling, and with anti-periodic boundary conditions.
@@ III.2.2 Single-site properties: non-Hermiticity score of number operator $\hat{n}$
Figure 5: Global characteristics of $\overline{{SC}^{\mathcal{F}}_{\mathrm{nh}}}$ for the non-Hermitian Hatano–Nelson model ($H_{\mathrm{nh}}^{\mathrm{HN}}$). We compute the total number of elements of $\overline{{SC}^{\hat{n}}_{\mathrm{nh}}}$ and $\overline{{SC}^{\mathcal{S}}_{\mathrm{nh}}}$ with the conditions $^{SC}^{\hat{n}}_{\mathrm{nh}}[k]\geq\mathcal{E}_{\mathrm{Th}}^{\hat{n}}$ and ${SC}^{\mathcal{S}}_{\mathrm{nh}}[k]\geq\mathcal{E}_{\mathrm{Th}}^{\mathcal{S}}$, respectively, where k=1,2,,Dk=1,2,...,D with $D$ being the Hilbert space dimension. These quantities, scaled by $D$, are shown in (a) and (b). We obtain a similar trend as in Fig. 4. Data for threshold values $\mathcal{E}_{\mathrm{T}h}^{\hat{n}}=\mathcal{E}_{\mathrm{T}h}^{\mathcal{S}}=0.1$. Similar to Figs. 1 and 4, the white circles correspond to the PT symmetry-breaking transition. The data is reported for $N=12$ sites, half-filling, and with anti-periodic boundary conditions.
[…]
@@
Unlike the case of $\hat{H}_{\mathrm{nh}}^{\mathrm{Bell}}$, here the spectrum of the Hamiltonian consists of a large number of eigenvectors. Therefore, to illustrate the non-Hermitian behavior across the spectrum of the model, in Fig. 5(a) we plot the quantity $\mathcal{G}^{\hat{n}}$, see Sec. II.0.2, choosing as threshold value $\mathcal{E}_{\mathrm{T}h}^{\hat{n}}=0.1$. The behavior is qualitatively similar to the one of $SC^{\hat{n}}_{\max}$, showing that in this case, the maximal score is representative of the behavior across the spectrum.
@@ IV Discussion
In this work, we have introduced a formalism that quantifies different notions of non-Hermiticity, which can emerge in any generic non-Hermitian system. In particular, as first quantifier, we have introduced the Hamiltonian non-Hermiticity, a measure that computes the normalized distance between a non-Hermitian Hamiltonian $\hat{H}_{\mathrm{nh}}$ and its hermitian conjugate $(\hat{H}_{\mathrm{nh}})^{\dagger}$. We have introduced a second quantifier to capture a notion of non-Hermiticity at the level of physically computable quantities related to the model. This quantifier has been motivated by noticing that in the non-Hermitiam paradigm one can prepare quantum ensembles of right or left eigenvectors, which often yield incompatible results with respect to single- and multi-site quantum properties. Based on this observation, we have introduced the non-Hermiticity score that characterized the difference between observables in the right and left eigenensemble. These quantifiers enable a comparative study of the degree of non-Hermiticity across different regions of parameter space, for static as well as dynamic properties.
@@
In summary, in our work, we have opened a new route to examining the quantum properties of any generic non-Hermitian system. It permits to quantify non-Hermiticity, not only along the traditional lines that solely focus on the Hamiltonian of the model but also with respect to its influence onto observable properties. There exist already several instances of models in the literature for which it is known that the left and right eigenvectors show drastically different physical behavior, e.g., in terms of localization [10, 33] or of topological properties [34, 35]. From this knowledge, one can immediately construct relevant non-Hermiticity scores that witness the non-trivial non-Hermitian behavior of the model. In the future, it will be interesting to apply our formalism to such and other physical systems to characterize exotic quantum many-body phases and identify potential non-Hermitian phase transitions.
An interesting complementary direction is to employ a bottom-up approach to construct non-Hermitian models with distinct entanglement properties—both bipartite and multipartite—as well as classical simulability features for their left and right eigenvectors. Such an approach could play a relevant role in preparing resourceful quantum states for advancing quantum technologies.